OpenStudy (anonymous):

how would you solve 1/3+1/2=1/X? Remember that / means a fraction.

OpenStudy (campbell_st):

well 1/3 + 1/2 = 5/6 divide the numerator and denominator by 5 to get 1/X

OpenStudy (campbell_st):

Oh... it will mean X is a decimal

OpenStudy (anonymous):

what were the steps you did to get 5/6?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well you need a common factor to add 2 fractions with different denominators... so 3 x 2 = 6 so the fractions you are adding will be equivalent fractions 1/3 = 2/6 and 1/2 = 3/6 so 2/6 + 3/6 = 5/6

OpenStudy (anonymous):

thank you so much.

OpenStudy (campbell_st):

so to look at your question properly you will have 5/6 = 1/x now you need to cross multiply to remove the denominators 5X = 6 divide both sides by 5 X = 6/5 or 1.2

OpenStudy (anonymous):

thank you!!